# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ Abstract
Traditional Visual Simultaneous Localization and Mapping (vSLAM) systems focus solely on static scene structures, overlooking dynamic elements in the environment. Although effective for accurate visual odometry in complex scenarios, these methods discard crucial information about moving objects. By incorporating this information into a Dynamic SLAM framework, the motion of dynamic entities can be estimated, enhancing navigation whilst ensuring accurate localization. However, the fundamental formulation of Dynamic SLAM remains an open challenge, with no consensus on the optimal approach for accurate motion estimation within a SLAM pipeline.
Therefore, we developed DynoSAM, an open-source framework for Dynamic SLAM that enables the efficient implementation, testing, and comparison of various Dynamic SLAM optimization formulations. DynoSAM integrates static and dynamic measurements into a unified optimization problem solved using factor graphs, simultaneously estimating camera poses, static scene, object motion or poses, and object structures. We evaluate DynoSAM across diverse simulated and real-world datasets, achieving state-of-the-art motion estimation in indoor and outdoor environments, with substantial improvements over existing systems. Additionally, we demonstrate DynoSAM’s utility in downstream applications, including 3D reconstruction of dynamic scenes and trajectory prediction, thereby showcasing potential for advancing dynamic object-aware SLAM systems. DynoSAM is open-sourced at https://github.com/ACFR-RPG/DynOSAM.
Keywords:
SLAM, Mapping, RGB-D Perception, Dynamic SLAM
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) has been a key area of research for over three decades [1]. Despite significant advancements, most SLAM systems are designed with the assumption of a predominantly static environment [2, 3, 4]. This limitation presents challenges in real-world scenarios, where dynamic objects are common and must be accounted for.
@@
Figure 1: DynoSAM is an open-source smoothing and mapping framework for Dynamic SLAM. The output of the system is shown in (a), which includes camera and object trajectories as well as the static and per-object dynamic map. (b) visualizes the feature-based front-end which performs multi-object tracking in addition to visual odometry. (c) shows dynamic map from the camera’s perspective, highlighting the estimated trajectory of each object and the tracked 3D points.
@@
Traditionally, SLAM systems classify sensor measurements associated with moving objects as outliers, excluding such data from the estimation process [5, 6], and in doing so, discard valuable information about dynamic elements. Incorporating these objects into the SLAM framework enables accurate modeling of dynamic environments [7, 8], which directly benefits navigation and task planning systems [9, 10] and enhances the overall robustness of SLAM [11, 12].
@@
Although learning-based methods have advanced robotic capabilities in dynamic environments, especially in autonomous vehicle applications with access to large datasets, there remains a critical need for SLAM solutions in scenarios where motion models are unknown and data is scarce. These challenges are particularly pronounced when operating in unstructured environments [13] or space exploration applications [14].
As a result, there is increasing interest in extending SLAM systems to incorporate observations of dynamic entities and estimate their motions [1, 15, 16, 17, 18, 19]. In this paper we refer to such systems as Dynamic SLAM.
Recent advances in Dynamic SLAM have explored methods such as multi-object visual odometry techniques [17, 20] and graph-based optimization approaches [6, 15, 21]. The latter methods aim to jointly estimate for the robot pose, static structure, and the motion or trajectory of rigid-body objects in the scene using both static and dynamic point observations. The literature presents diverse approaches to formulating the Dynamic SLAM problem, each characterized by a distinct set of variables and objective functions, leading to different underlying graph structures. A recent study [18] compared object-centric approaches, where dynamic points or motions are represented in an object’s body frame, with world-centric approaches, where points and motions are modeled in the world frame. The findings demonstrated that the choice of variables and formulation has a substantial impact on the robustness, accuracy, and efficiency of SLAM systems.
Despite the growing interest in Dynamic SLAM research, there remains limited consensus on the best formulation to achieve accurate motion estimation. This motivated us to develop DynoSAM, a framework that provides a structured approach for implementing and evaluating Dynamic SLAM solutions. The proposed framework implements, evaluates and compares both existing approaches and new formulations for object motion and pose estimation, including those based on a world-centric [11, 15, 18] paradigm.
To rigorously evaluate these formulations, we have integrated them into a full Dynamic SLAM system that uses visual (stereo/RGB-D) sensor data as input, where Fig. 1 presents an example of our system estimating complex dynamic object motions in the Oxford Multimotion Dataset (OMD) [22]. Our framework uniquely models the motion of each object as a rigid-body pose change [23], offering flexibility in the parametrization of our estimation problem. We assess the accuracy and performance of our system using a variety of indoor and outdoor datasets.
DynoSAM is compared against other Dynamic SLAM pipelines [17, 15, 24, 21], where we achieve state-of-the-art accuracy in object motion and pose estimation. Furthermore, we demonstrate how our framework can be used to better inform downstream navigation applications by integrating DynoSAM with dynamic object reconstruction [8] and trajectory prediction methods.
II Contributions
The specific contributions of this paper are as follows:
  • A framework for Dynamic SLAM that provides a structured approach for implementing, testing and evaluating graph-based solutions. Our framework establishes common terminologies and metrics for evaluating Dynamic SLAM pipelines, facilitating more robust conclusions and analysis. Within our framework we present a customizable recipe of formulations that serve as the theoretical foundation for a practical Dynamic SLAM system. We present a discussion and analysis for each method.
  • A complete Dynamic Visual SLAM pipeline that incorporates all presented formulations in this paper. We evaluate the accuracy of our system on a variety of datasets.
  • A novel formulation for Dynamic SLAM that directly parametrizes the pose of each object in the scene, whilst ensuring rigid-body kinematics are maintained.
  • An exhaustive evaluation of our framework and proposed methodologies across various datasets, demonstrating state-of-the-art performance. To the best of our knowledge, this paper provides one of the most extensive evaluations of Dynamic SLAM formulations in the existing literature.
  • An implementation of a sliding-window optimization approach to bound the computation. We analyze the impact on overall accuracy of the sliding-window compared to a full-batch approach.
  • Integration of our Dynamic SLAM pipeline with downstream tasks highly relevant for navigation, such as dynamic object reconstruction [8] and trajectory prediction.
@@ II Contributions
As part of our contribution we open-source our C++ implementation of DynoSAM, which is integrated with the Robotic Operating System 2 (ROS2) [25] and can be run with offline datasets or online data. Additionally, we provide data logging and serialization tools for each module to facilitate evaluation and debugging as well as an accompanying automated evaluation suite that has been used to generate all results in this paper. To facilitate modularity and parallelization of key components, DynoSAM’s architecture is inspired by Kimera [26].
III Related Work
@@
This section provides an overview on how SLAM systems have been addressing the challenges of dynamic environments over the years. We discuss SLAM systems that aim to improve their robustness by removing or ignoring dynamic objects, as well as state-of-the-art Dynamic SLAM systems that estimates object poses and motions.
@@
In order to handle the dynamic entities in the environment, conventional SLAM systems such as ORB-SLAM 3 [4] and DynaVINS [27] detect any moving observations as outliers and reject them to create a global map that only contains static structures using methods such as RANSAC [28], point correlations [29] and robust bundle adjustment formulation [27, 30]. Deep learning methods have also been recently used to semantically understand the scene, and detect and remove dynamic objects [6, 31, 32]. By taking an active approach in removing dynamic objects, these methods provide robust and accurate camera pose estimation in dynamic environments. However, any relevant information about the motion of the objects is discarded. Khronos [33] proposes a spatio-temporal SLAM formulation that attempts to unify short-term changes (such as dynamic objects) with long-term scene changes. They extract moving objects from their representation using a geometric motion detection method [34] and are able to construct a comprehensive map in complex, changing environments. However, they do not utilize any estimate of object motion and rely on a dense reconstruction pipeline to detect and track objects in the scene.
@@
To represent the kinematic information of these objects, Dynamic SLAM methods incorporate measurements of dynamic entities in the SLAM formulation in addition to the static ones. These systems first segment dynamic observations from the static background using information such as kinematics [24, 17] and semantics [15, 21, 35], before optimizing the pose or motion of these objects jointly with the camera/robot poses and the map of the environment. Barsan et al[36] leveraged semantic and kinematic segmentation to detect dynamic objects, tracked their motions using sparse scene flow and map them separately from the static scene, but they only handled frame-to-frame estimation without any joint optimization [36].
Within the Dynamic SLAM formulations that explicitly represent the object, the most common and intuitive approach is to directly estimate the object’s pose [21, 35, 37, 38]. Assuming each object is a rigid body, observed points on these objects are static in a body-fixed local object reference frame. By defining the poses of objects and in turn their local reference frames, there is an immediate advantage that each object point can be expressed in the object reference frame with a single variable in the optimization problem, reducing the overall number of variables in the system. Huang et al[24, 35] cluster point observations on moving objects in a scene based on temporal rigidity to identify underlying rigid bodies. The resulting clusters of points defines the pose of each object and the poses of the camera and each dynamic object is solved through a sliding-window optimization. Their method, although able to extract dynamic objects from the scene, demonstrates poor object motion accuracy. Alternatively, several methods model objects using simple geometric primitives such as cuboids [39, 40, 41, 42] and ellipsoids [43] so as to define their poses. Recently, DynaQuadric [19] has proposed a joint optimization framework that reconstructs dynamic objects as quadrics. This unique representation allows the per-object scale to be explicitly defined and they demonstrate highly accurate camera localization. However, they evaluate only their object pose errors against other quadric-based methods and neglect object motion evaluations.
@@
Several neural network-based methods have also been proposed in recent years [44, 45, 46, 47] to estimate the scene flow of 3D points or voxels. These systems focus on estimating the motions of individual points in the form scene flow vectors. While they produce impressive results, these systems do not recover the motions or poses of dynamic objects, and are therefore not further discussed in this paper.
@@
When focusing on specific object types, BodySLAM++ [48] and DSP-SLAM [49] leverage learned object shape prior models in their formulation and achieve accurate object pose estimation. While their results demonstrate the benefit of integrating learned priors into a graph-based optimization framework, BodySLAM++ is restricted to only estimating human poses and DSP-SLAM provides no understanding on object dynamics. TwistSLAM [41] utilizes mechanical joint constraints to restrict the degrees of freedom for object motion estimation. While they demonstrate accurate pose estimation under constrained motion, their results suggest limitations in estimating the full $\mathrm{SE}(3)$ motion of objects, especially for objects with more complex motion patterns.
@@
By comparison, other methods seek to estimate the full $\mathrm{SE}(3)$ motion of each object without the addition of semantic-specific constraints. Among these solutions, DynaSLAM II [21] takes an object-centric approach to motion estimation and reports the most accurate egomotion when compared with other approaches. By comparison, their experimental results demonstrate poor object motion estimation and the authors consider their use of sparse features to be the main reason behind such performance. However, our previous work [18] shows that the formulation used by [21] does not explicitly model the rigid-body kinematics, resulting in poor estimation accuracy even with a dense object representation.
@@
An alternative formulation is to estimate object motions directly in a known reference frame, such as a camera frame that moves with a sliding window  [17, 20, 22], or a well-defined reference frame like the world frame [18, 15, 50], which commonly coincides with the first camera/robot pose. MVO [17] represents the dynamic points in the camera frame, but models object motions in the object frame. They define object poses using point observations when each object is first observed, similar to the method of Bescos et al[21].
@@
Chirikjian et al[23] demonstrated that a $\mathrm{SE}(3)$ motion can be expressed in any reference frame, and based this concept, VDO-SLAM [15, 11, 51] proposes a factor graph formulation to represent and estimate rigid-body object motions in the world frame without the need to estimate object poses. AirDOS [50] employs the same rigid-body motion model and extends it to articulated objects, e.g. humans, by combining it with a learning-based human pose estimator [52] to estimate for the motion of each segment of an articulated object. Interestingly, they quantitatively demonstrate direct improvement in their camera pose estimation due to the incorporation of dynamic objects. While the study of non-rigid body motion outside the scope of this paper, we find such a strategy inspiring for our future works. Our previous work [18] investigated the motion model employed by [15, 50] by comparing world and object-centric formulations and demonstrated that the world-centric approach better encodes rigid-body kinematic models, leading to more accurate motion estimation. Additionally, the world-centric approach exhibits more efficient convergence behavior and often requires shorter computation time despite requiring more variables. Development of [18] also highlighted practical challenges due to the limited availability of open-source Dynamic SLAM tools and systems which hindered rigorous comparison and evaluation. These limitations motivated the development of DynoSAM as a flexible and comprehensive framework for Dynamic SLAM.
IV Background
IV-A Rigid-Body Motion

### v3

@@ Abstract
Traditional Visual Simultaneous Localization and Mapping systems focus solely on static scene structures, overlooking dynamic elements in the environment. Although effective for accurate visual odometry in complex scenarios, these methods discard crucial information about moving objects. By incorporating this information into a Dynamic SLAM framework, the motion of dynamic entities can be estimated, enhancing navigation whilst ensuring accurate localization. However, the fundamental formulation of Dynamic SLAM remains an open challenge, with no consensus on the optimal approach for accurate motion estimation within a SLAM pipeline.
Therefore, we developed DynoSAM, an open-source framework for Dynamic Objects SLAM that enables the efficient implementation, testing, and comparison of various Dynamic SLAM optimization formulations. We further propose a novel formulation that encodes rigid-body motion model in object pose estimation as well as an error metric agnostic to object frame definition. DynoSAM integrates static and dynamic measurements into a unified optimization problem solved using factor graphs, simultaneously estimating camera poses, static scene, object motion or poses, and object structures. We evaluate DynoSAM across diverse simulated and real-world datasets, achieving state-of-the-art motion estimation in indoor and outdoor environments, with substantial improvements over existing systems. Additionally, we demonstrate DynoSAM’s contributions to downstream applications, including 3D reconstruction of dynamic scenes and trajectory prediction, thereby showcasing potential for advancing dynamic object-aware SLAM systems. Code is open-sourced at https://github.com/ACFR-RPG/DynOSAM
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) has been a key research area for over three decades [Rosen2021annurev]. Despite significant advancements, most SLAM systems are designed with the assumption of a predominantly static environment [Newcombe2011ismar, mur2017orb, Campos2021tro]. This limitation presents challenges in real-world scenarios, as dynamic objects are common and must be accounted for.
@@
Figure 1: DynoSAM is an open-source smoothing and mapping framework for Dynamic SLAM. (a) presents the system output, which includes camera and object trajectories, as well as the static and per-object dynamic map. (b) visualizes the feature-based front-end which performs multi-object tracking in addition to visual odometry. (c) shows the dynamic map from the camera’s perspective, highlighting the estimated trajectory of each object and the tracked 3D points.
@@
Traditionally, SLAM systems classify sensor measurements associated with moving objects as outliers, excluding such data from the estimation process [Zhao08icra, Bescos2018ral], and in doing so, discard valuable information about dynamic elements. Incorporating these objects into the SLAM framework enables accurate modeling of dynamic environments [Newcombe2015CVPR_dynamicfusion, Wang2025icra_dynorecon], which directly benefits navigation and task planning systems [Finean2021iros, Hermann2014] and enhances the overall robustness of SLAM [Henein20icra, Wang07ijrr].
@@
Although learning-based methods have advanced robotic capabilities in dynamic environments, especially in autonomous vehicle applications with access to large datasets, there remains a critical need for SLAM solutions in scenarios where motion models are unknown and data is scarce. These challenges are particularly pronounced when operating in unstructured environments [deeb2021quasistatic] or space exploration applications [setterfield2018mapping].
As a result, there is increasing interest in extending SLAM systems to incorporate observations of dynamic entities and estimate their motions [Rosen2021annurev, zhang2020vdoslam, Liu2024ral_dynameshslam, judd2024ijrr_mvo, morris2024icra, Tian2024its_dynaquadric]. In this paper we define the term Dynamic SLAM as SLAM solutions that optimize for object motions and/or poses in addition to robot trajectory and the static structure, to distinguish from those that exclude objects to be robust in dynamic environments [Bescos2018ral, Song2022ral_dynavins, song2024dynavins++]. This paper focuses on studying purely Visual SLAM (vSLAM) formulations, though we have further discussed IMU-integrated experiments and results.
There exists limited consensus in the community for how to best formulate the Dynamic SLAM problem. The seminal work of Henein et al[Henein20icra] showed that the estimation problem can be designed around estimating for the full $\mathrm{SE}(3)$ object motion. These motions are jointly optimized with the camera poses, and the static and dynamic structures, from which the per-object trajectory and velocities can be recovered. Their following work, VDO-SLAM [zhang2020vdoslam], extended this formulation into a complete pipeline, but was never formally published. The works of Judd et al[judd2024ijrr_mvo, Judd18iros] similarly prioritize object motion, using this information to segment moving objects and track objects through occlusion. Comparatively, many systems directly estimate for object poses [bescos2021ral, Huang2019iccv, huang2020cvpr], typically using a constant-motion model fully constrain the estimation. However, downstream applications such as reconstruction [Wang2025icra_dynorecon], navigation [Phillips2011icra_sipp] and prediction[salzmann2020trajectron++] often require information on both object pose and motion.
Our prior work [morris2024icra] identified a further point of comparison between Dynamic SLAM formulations. Instead of only considering the estimation of either object pose or motion, this study termed and compared object-centric formulations, where dynamic points and motions are represented in an object’s body frame, with world-centric formulations, which represent the variables in the world-frame. The insights from our work demonstrated that the choice of formulation plays a vital role in determining the behaviour of convergence and the resulting estimation accuracy. Using the world-centric motion model [Henein20icra, zhang2020vdoslam, morris2024icra] this study showed that explicit modelling of the object’s rigid-body kinematics is essential for accurate results, especially when direct estimation of the object pose is required.
The inconsistencies in results and the variety of approaches, coupled with the communities growing interest in Dynamic SLAM motivated us to develop Dynamic Object Smoothing And Mapping, DynoSAM, a framework that allows for easy and modular prototyping, implementation and evaluation of Dynamic SLAM formulations. The suffix ‘-SAM’ is inherited from the GTSAM [gtsam] library which our framework is built upon, thereby allowing highly customisable estimation problems to be designed and implemented.
In this work, we present a detailed theoretical foundation for Dynamic SLAM based on a world-centric model for object motion. Furthermore, to support downstream tasks requiring object poses, we propose a novel formulation that explicitly uses the rigid-body kinematic model for accurate, direct pose estimation. To rigorously evaluate these formulations, we have integrated them into a full Dynamic vSLAM system; Fig. 1 presents an example output of our system estimating complex dynamic object motions. We assess the accuracy and performance of our system using a variety of indoor and outdoor datasets. To standardise and facilitate valid comparisons between systems, we additionally propose a new error metric for evaluating object motion. Compared with other dynamic SLAM pipelines [zhang2020vdoslam, judd2024ijrr_mvo, Huang2019iccv, bescos2021ral], we achieve state-of-the-art accuracy in the estimation of motion and pose of objects. We demonstrate that our camera pose estimation is robust in highly dynamic environments and compare with AirDOS [Qiu2022icra_airdos], DynaVINS [Song2022ral_dynavins] and ORB-SLAM3 [Campos2021tro]. Finally, we demonstrate that our formulations can directly inform downstream navigation tasks by integrating our framework with dynamic object reconstruction [Wang2025icra_dynorecon] and trajectory prediction methods.
In summary, the contribution of this paper is threefold:
  • Theoretical Foundations for Object Motion Estimation. We present a detailed theoretical formulation for representing and integrating observations of rigid bodies in motion within Dynamic SLAM. We offer a set of formulations that form the theoretical basis for practical system development, including a novel Dynamic SLAM formulation that directly parametrizes each object’s pose in the scene and explicitly encodes rigid-body kinematics.
  • A Modular and Practical Dynamic vSLAM System. We implement a complete Dynamic Visual SLAM system that integrates the various formulations discussed, supporting both batch and sliding-window optimizations. In addition, we provide a practical demonstration of our Dynamic SLAM pipeline with downstream tasks highly relevant for navigation, such as dynamic object reconstruction [Wang2025icra_dynorecon] and trajectory prediction. The framework is modular and extensible, allowing researchers to easily test and benchmark new formulations within a consistent evaluation setup.
  • Extensive Benchmarking and Novel Evaluation Metric. We introduce an object motion evaluation metric that is agnostic to object reference frame definitions. Our framework is exhaustively evaluated on multiple public datasets, achieving state-of-the-art results. To our knowledge, this work provides one of the most comprehensive evaluations of Dynamic SLAM formulations to date.
@@
As part of our contribution we open-source our C++ implementation of DynoSAM, which is integrated with the Robotic Operating System 2 (ROS2) [Thomas2014Ros2]. We provide data logging and serialization tools for each module to facilitate evaluation and debugging, as well as an accompanying automated evaluation suite that has been used to generate all results in this paper. To facilitate modularity and parallelization of key components, DynoSAM’s implementation is inspired by Kimera [Rosinol20icra].
II Related Work
@@
This section provides an overview on how SLAM systems address the challenges of dynamic environments over the years. We focus on state-of-the-art Dynamic SLAM systems that estimate object poses and motions, but we also include those systems that aim to improve robustness by removing or ignoring dynamic objects.
II-A SLAM in Dynamic Environments
@@
In order to handle the dynamic entities in the environment, conventional SLAM systems such as ORB-SLAM 3 [Campos2021tro] and DynaVINS [Song2022ral_dynavins] detect any moving observations as outliers and reject them to create a global map that only contains static structures using methods such as RANSAC [fischler1981cacm], point correlations [Dai2022pami] and robust bundle adjustment formulation [Song2022ral_dynavins, song2024dynavins++]. Deep learning methods have also been recently used to semantically understand the scene, and detect and remove dynamic objects [Bescos2018ral, hachiuma2019detectfusion, zhang2020flowfusion]. By taking an active approach in removing dynamic objects, these methods provide robust and accurate camera pose estimation in dynamic environments. However, any relevant information about the motion of the objects is discarded. Khronos [Schmid2024rss_Khronos] operate in dynamic environments by proposing a spatio-temporal SLAM formulation that unifies short-term changes, e.g. from dynamic objects, with long-term scene changes, and extracts moving objects from their representation using a geometric motion detection method [Schmid2023dynablox]. However, they do not utilize any estimate of object motion and rely on a dense reconstruction pipeline to detect and track objects in the scene.
To improve the robustness of camera pose estimation in dynamic environments, many of these methods either implicitly or explicitly identify and segment out moving objects in the scene. While motion detection and segmentation is highly relevant for Dynamic SLAM, we highlight that our work is primarily interested in the estimation of object motions and poses, in addition to segmenting motions.
II-B Dynamic SLAM
@@
To represent the kinematic information of these objects, Dynamic SLAM methods incorporate measurements of dynamic entities in the SLAM formulation in addition to the static ones. These systems first segment dynamic observations from the static background using information such as kinematics [judd2024ijrr_mvo, Huang2019iccv] and semantics [zhang2020vdoslam, bescos2021ral, huang2020cvpr], before solving for the pose or motion of these objects in addition to the camera/robot poses and the map of the environment either via joint optimization [morris2024icra, bescos2021ral] or separate estimation, such as frame-to-frame object motion tracking using sparse scene flow [Barsan2018icra].
@@
Several neural network-based methods have also been proposed in recent years [zhang2024arxiv_deflow, zhang2024arxiv_seflow, Wang2023pami, Li2023iccv] to estimate the scene flow of 3D points or voxels. These systems focus on estimating the motions of individual points in the form scene flow vectors. While they produce impressive results, these systems do not recover the motions or poses of dynamic objects, and are therefore not further discussed in this paper.
II-B1 Object Pose Representation
Within the Dynamic SLAM formulations that explicitly model the object, the most common and intuitive representation is to directly estimate the object’s pose [bescos2021ral, huang2020cvpr, Wang2007ijrr_SLAMMOT, Ballester2021icra]. Assuming each object is a rigid body, observed points on these objects are static in a body-fixed local object reference frame. By defining the poses of objects and in turn their local frames, there is an immediate advantage that each object point can be expressed in the object reference frame with a single variable in the optimization problem, reducing the overall number of variables in the system. Huang et al[Huang2019iccv, huang2020cvpr] cluster point observations on moving objects based on temporal rigidity to identify underlying rigid bodies. Their method, although able to extract dynamic objects from the scene, demonstrates poor object motion accuracy. Alternatively, several methods model objects using simple geometric primitives such as cuboids [Yang2019tro, li2018eccv, gonzalez2022twistslam, gonzalez2023twistslam++] and ellipsoids [nicholson18cvpr] to define their poses. DynaQuadric [Tian2024its_dynaquadric] has proposed a joint optimization framework that models dynamic objects as quadrics, explicitly defining the per-object scale and demonstrating highly accurate camera localization. However, they evaluate only their object pose errors against other quadric-based methods and neglect object motion evaluations.
@@ II-B1 Object Pose Representation
When focusing on specific object types, BodySLAM++ [Henning2023iros_bodyslam++] and DSP-SLAM [Wang20213dv_dspslam] leverage learned object shape prior models and achieve accurate object pose estimation. Their results demonstrate the benefit of integrating learned priors into a graph-based optimization framework, but BodySLAM++ is restricted to only estimating human poses and DSP-SLAM provides no understanding on object dynamics. Assuming strong knowledge on object motion, TwistSLAM [gonzalez2022twistslam] uses mechanical joint constraints to restrict the degrees of freedom for object motion estimation, and estimates accurate object poses under constrained motion. Their results however suggest limitations in estimating complex $\mathrm{SE}(3)$ object motions.
@@
In comparison, other methods seek to estimate the full $\mathrm{SE}(3)$ motion of each object without the addition of semantic-specific constraints. Among these solutions, DynaSLAM II [bescos2021ral] takes an object-centric approach to pose and velocity estimation and reports the most accurate egomotion when compared with other approaches. By comparison, DynaSLAM II presents poor object motion estimation, and the authors consider their use of sparse features as the main reason behind such performance. However, our prior work [morris2024icra] shows that the formulation used by Bescos et al[bescos2021ral] does not explicitly model the rigid-body kinematics, resulting in poor estimation accuracy even with a dense object representation.
II-B2 Object Motion Representation
@@
An alternative formulation is to estimate object motions directly in a known reference frame, such as a camera frame that moves with a sliding window  [judd2024ijrr_mvo, Judd18iros, Judd19ral], or a well-defined reference frame like the world frame [zhang2020vdoslam, morris2024icra, Qiu2022icra_airdos], which commonly coincides with the first camera/robot pose. MVO [judd2024ijrr_mvo] represents the dynamic points in the camera frame, but models object motions in the object frame. They define object poses using point observations when each object is first observed, similar to the method of Bescos et al[bescos2021ral].
@@
Chirikjian et al[Chirikjian17idetc] showed that a $\mathrm{SE}(3)$ motion can be expressed in any reference frame, and based on this concept, VDO-SLAM [Henein20icra, zhang2020vdoslam, zhang20iros] proposes a factor graph formulation to represent and estimate rigid-body object motions in the world frame without the need to estimate object poses. AirDOS [Qiu2022icra_airdos] employs the same rigid-body motion model, extends it to articulated objects, e.g. humans, via a learning-based human pose estimator [Fang2017iccv_rmpe], and estimates the motion of each segment of an articulated object. They quantitatively demonstrate direct improvement in camera pose estimation due to the incorporation of dynamic objects. While the study of non-rigid body motion outside the scope of this paper, we find such a strategy inspiring for our future works. The motion models employed by [zhang2020vdoslam, Qiu2022icra_airdos] are interrogated in [morris2024icra] which showed, by comparing world and object-centric formulations, that a world-centric approach better encodes rigid-body kinematic models, leading to more accurate motion estimation, more efficient convergence behaviour and often shorter computation time, despite requiring more variables.
A comprehensive review of the relevant literature highlights practical challenges, including the scarcity of open-source Dynamic SLAM libraries, systems, and datasets, which prevent rigorous comparison and evaluation. Inconsistent definitions of object reference frames further complicate pose and motion assessment. These limitations motivated the development of DynoSAM, a flexible and comprehensive framework for Dynamic SLAM with frame-agnostic error metrics for evaluation.
III Background
III-A Rigid-Body Motion